OpenStudy (precal):

Give the location of the vertical asymptote(s) if they exist, and state the functions domain. I believe the textbook's solution may be incorrect

OpenStudy (precal):

\[h(x)=\frac{x^2-1}{2x^2+3x-5}\]

OpenStudy (anonymous):

\[h(x)=\frac{(x-1)(x+1)}{(x-1)(2x+5)}\] it has 1 vertical asymptote at x=-5/2

OpenStudy (precal):

I factored them and got \[h(x)= \frac{(x+1)(x-1)}{(x+\frac{5}{2})(x-1)}\]

OpenStudy (precal):

I wrote that the same v.a value but my book states x=1 also. I believe that is a hole. For the domain